OpenStudy (anonymous):

Pete Moran bought a Dell Laptop Computer priced at $699. He put down 30 percent. The amount of down payment is:

OpenStudy (anonymous):

"Pete Moran bought a Dell Laptop Computer " Big mistake...:-)

OpenStudy (anonymous):

\[\frac{30}{100}=\frac{x}{699}]\]

OpenStudy (anonymous):

lol estudier!

OpenStudy (anonymous):

cross multiply and solve

OpenStudy (anonymous):

30*699=100x

OpenStudy (anonymous):

209700=100x

OpenStudy (anonymous):

2097=x

OpenStudy (anonymous):

$2,097

OpenStudy (anonymous):

$209.7 .3*699 = $209.7